A purse contains $6.50 in quarters and dimes. There are 5 more quarters than dimes.Find the number of quarters and dimes
Answer by stanbon(75887)

A purse contains $6.50 in quarters and dimes. There are 5 more quarters than dimes.Find the number of quarters and dimes
----
Value Equation::: 10d + 25q = 650 cents
Quantity Equatin: q = d+5
----
Substitute for "q" and solve for "d":
10d + 25(d+5) = 650
---
10d + 25d + 125 = 650
35d = 525
d = 15 (# of dimes)
----
Solve for "q":
q = d+5
q = 15+5 = 20 (# of quarters)
================================
Cheers,
Stan H.
=========